OpenStudy (anonymous):

can anyone check my answer? [5x/(x+1)]-[(1+4x)/(x+1)] i got [(x-1)/(x+1)]

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

OpenStudy (unklerhaukus):

but you can simplify a bit further

OpenStudy (anonymous):

i can? -1?

OpenStudy (unklerhaukus):

[(x-1)/(x+1)] = [(x+1-2)/(x+1)] = [(x+1)/(x+1)-(2)/(x+1)] =

OpenStudy (anonymous):

that lost me

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (unklerhaukus):

now look to cancel